In the laboratory a student finds that it takes 56.4 Joules to increase the temperature of 11.0 grams of solid chromium from 23.7 to 35.9 degrees Celsius. The specific heat of chromium calculated from her data is ____ j/gC
1
Expert's answer
2020-06-29T07:45:01-0400

In the laboratory a student finds that it takes 56.4 Joules to increase the temperature of 11.0 grams of solid chromium from 23.7 to 35.9 degrees Celsius. The specific heat of chromium calculated from her data is 0.420 J/g°C


C = Q / (m×T )         C= Specific heat, Energy(Q)= 56.4 J, Mass(m)= 11 g,

C = 56.4 /(11.0×12.2) J/(g×°C)                Change of temperature (T) = (35.9-23.7)°C

0.420 J/(g×°C) (Answer)                                                                =12.2°C